OpenStudy (anonymous):

5 - x/3 = 3(x + 1)/7

OpenStudy (anonymous):

I had the Idea to multiply the denominator

OpenStudy (paki):

yeah so what will you multiply... term...?

OpenStudy (anonymous):

3

OpenStudy (paki):

"3" with what...?

OpenStudy (anonymous):

after closing my ( ) I should have 5 - x/3 = 3x + 3/7

OpenStudy (paki):

ok see.... you will multiply each side of the equation by (21) so that the equation may be simplified and free from divisions....

OpenStudy (anonymous):

I thought it was 3(5) - 3(x)/3 = 3(3x = 3(3x + 3)/7

OpenStudy (paki):

just do this, what i told, and show your work here please...

OpenStudy (paki):

no... there are two values in the denominator... (3 and 7)... so it makes 3*7=21.... so we will multiply both sides of the equation by "21"... ok....

OpenStudy (anonymous):

ok see 5 - x = 3x = 3 but to where am I multiplying 21

OpenStudy (paki):

no... not like this... i am showing to you... ok...

OpenStudy (anonymous):

3x= 8

OpenStudy (paki):

5 - x/3 = 3x + 3/7... ok now multiply each term by "21"... 21(5) -21 (x/3) = 21(3x+3/7).... 105 - 7x = 9x+9... got this....

OpenStudy (anonymous):

16x = 114

OpenStudy (anonymous):

57/8 = x

OpenStudy (paki):

no... it will be.... 16x=96.... x= 6...

OpenStudy (anonymous):

oh because you have to -9 got it Idk why I got confused

OpenStudy (paki):

hmm... yup... these types of questions are very easy... just need little concentration... ok...

OpenStudy (anonymous):

yea

OpenStudy (paki):

:)